OpenStudy (anonymous):

I had to find the surface area of a cylinder and I have to write the answer in terms of pi how do I do that when I got 113.04? Do I just add pi to it? ex 113.04πyd62?

ganeshie8 (ganeshie8):

Nope.

ganeshie8 (ganeshie8):

what are the dimensions of cylinder ?

OpenStudy (anonymous):

r=2 yds h=7yds ?

OpenStudy (anonymous):

What is the formula for surface area of a cylinder

ganeshie8 (ganeshie8):

whats the formula for surface area ?

OpenStudy (anonymous):

is it yards

OpenStudy (anonymous):

yes yards and A=2πrh+2πr^2

ganeshie8 (ganeshie8):

factor out \(\pi\) and evaluate remaining stuff

ganeshie8 (ganeshie8):

A=2πrh+2πr^2 = π(2rh + 2r^2) = π(??? )

OpenStudy (anonymous):

this is what I got =π(28+8) = π(36) =113.04 or am I wrong?

ganeshie8 (ganeshie8):

you're absolutely correct ! but you did more than asked

ganeshie8 (ganeshie8):

this is what I got =π(28+8) = π(36) ********************************* =113.04 or am I wrong?

ganeshie8 (ganeshie8):

you must stop there itself, cuz you're asked to leave the answer in terms of π

OpenStudy (anonymous):

give the answer in terms of pi

OpenStudy (anonymous):

oh so it would just be 36π? or as you stated π36?

ganeshie8 (ganeshie8):

36π looks more neat

OpenStudy (anonymous):

It also asks for the unit whiihc is yards so I would add yds62 too? Or leave as is?

OpenStudy (anonymous):

Here 36 is a co-efficient of pi We always write co-efficients first

OpenStudy (anonymous):

yd^2*

OpenStudy (anonymous):

35 pi yd^2

OpenStudy (anonymous):

Alright thank you both so much for helping and explaining this to me.

OpenStudy (anonymous):

No problem and wc

ganeshie8 (ganeshie8):

np :) btw with units, its `36π yd^2`